OpenStudy (anonymous):

would \[2\sqrt{8}\]+\[3\sqrt{8}\]=\[5\sqrt{8}\]?

hartnn (hartnn):

yes, thats true once you factor out sqrt 8

hartnn (hartnn):

and that can be simplified further

hartnn (hartnn):

any doubts?

OpenStudy (anonymous):

umm i cant see how it can be simplified?

OpenStudy (anonymous):

You can still simplify your final answer: \[5\sqrt{8} = ?\]

hartnn (hartnn):

can you factor 8?

OpenStudy (anonymous):

8 2 & 4 2 & 2 & 2 ?

hartnn (hartnn):

good, so, \(\large \sqrt8=\sqrt{4\times 2}=\sqrt4 \times \sqrt2=....?\) and you know whats the square root of 4 ?

OpenStudy (anonymous):

\[2\sqrt{2}\]?

hartnn (hartnn):

absolutely correct :) so 5 sqrt 8 will be = 10 sqrt 2 got it ?

OpenStudy (anonymous):

so you times the 5 by the 2 outside, but not the 2 inside?

OpenStudy (anonymous):

2 by 5**

hartnn (hartnn):

yes, the \(\sqrt 2\) remains as is . \(5\sqrt 8 = 5\times 2\sqrt 2 =10 \sqrt2\)

OpenStudy (anonymous):

oh.. i should have written it out!

OpenStudy (anonymous):

thanks very much!

hartnn (hartnn):

welcomes very much ^_^